if 6 times a number is added to the square of the number, the result is 16. find the number(s)

Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39617)

What help do you need? The description is very plain.

6n+n^2=16
n(6+n)=16
The number is 2.

Answer by ikleyn(52778)

.
if 6 times a number is added to the square of the number, the result is 16. find the number(s)
~~~~~~~~~~~~~~~~~

n^2 + 6n = 16.


n^2 + 6n - 16 = 0.


(n-2)*(n+8) = 0.


The problem has 2 (two, TWO) solutions. These solutions are the numbers 2 and -8.    ANSWER

Solved.